OpenStudy (anonymous):

ax+2xy=14 solve for y

Nnesha (nnesha):

subtract ax both side

OpenStudy (anonymous):

y=-1/2a+7/x

OpenStudy (anonymous):

2xy=14-ax

Nnesha (nnesha):

now divide both side by 2x

Nnesha (nnesha):

let me know what you get @mathman13

OpenStudy (anonymous):

y=14-ax over 2x

OpenStudy (anonymous):

now simplify the right side

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

turn it into 14/2x - ax/2x

OpenStudy (anonymous):

then simplify

Nnesha (nnesha):

 divide 14 by 2

OpenStudy (anonymous):

ok then what